Circular No. 6827

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

4U 1630−47

R. M. Hjellming, National Radio Astronomy Observatory; and E. Kuulkers, Oxford University, report: “Using the VLA we have detected a possible variable radio counterpart to the recurring x-ray transient 4U 1630−47 (IAUC 6822, 6823). On Feb. 10.6, 12.6, 14.6, 21.6 and 25.5 UT a radio source at a position of α = 16^{h}34^{m}01^{s}.61 ± 0^{s}.05, δ = −47^{o}23′32′′ ± 2′′ (equinox 2000.0) had 4.9-GHz fluxes of < 1, 1.1 ± 0.3, 1.4 ± 0.3, 1.3 ± 0.3 and 0.5 ± 0.3 mJy, respectively. These flux levels, together with the upper limits of 0.7 mJy at 8.6 GHz from Kuulkers et al. (IAUC 6822) on Feb. 9.1 and 9.8, may indicate the variability needed to identify this radio candidate as the counterpart of the 1998 outburst in 4U 1630−47. The position of the radio source is consistent with the Einstein HRI x-ray position of α = 16^{h}34^{m}01^{s}.3, δ = −47^{o}23′34′′.3 (equinox 2000.0) reported by Parmar et al. (1997, A.Ap. 319, 855).”

M. Buxton and R. Sood, Australian Defence Force Academy; D. Rayner, University of Tasmania; T. Tzioumis, Australia Telescope National Facility; E. Kuulkers, Oxford University; S. Dieters, University of Alabama; and P. Durouchoux, Centre d’Etudes de Saclay, report: “Following earlier radio observations with the ATCA on Feb. 9.1 and 9.8 UT that set an upper limit of 0.7 mJy (IAUC 6822), a positive flux has been detected from within the x-ray error circle of 4U 1630−47 during Feb. 17.57–17.92. The coordinates of the radio source are α = 16^{h}34^{m}01^{s}.6, δ = −47^{o}23′34′′ (equinox 2000.0). The observed flux and polarization were: at 4800 MHz (6 cm), 2.6 ± 0.3 mJy, 26 percent; at 8640 MHz (3 cm), 1.9 ± 0.1 mJy, 28 percent.”

COMET 80P/PETERS-HARTLEY CCD observations by C. Veillet with the Canada-France-Hawaii Tele-

scope: Feb. 16.47 UT, R = 20.4; 17.51, 20.4; 18.58, 20.4.

COMET 43P/WOLF-HARRINGTON Total visual magnitude estimates: 1997 Nov. 12.19 UT, 12.7 (R. J.

Bouma, Groningen, The Netherlands, 0.25-m reflector); Dec. 2.79, 12.8 (A. Pearce, Nedlands, Western Australia, 0.20-m reflector); 7.11, 12.3 (T. Karhula, Virsbo, Sweden, 0.44-m reflector); 27.61, 12.8 (N. Biver, Oahu, HI, 0.26-m reflector); 1998 Jan. 5.60 (D. A. J. Seargent, Cowra, New South Wales, 0.25-m reflector); 20.55 13.3: (M. Mattiazzo, Wallaroo, South Australia, 0.20-m reflector); 29.96, 12.5: (K. Hornoch, Lelekovice, Czech Republic, 0.35-m reflector); Feb. 17.53, 13.0: (Mattiazzo).

1998 February 28  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Brian G. Marsden